The manager of a restaurant determined that the odds against a customer ordering dessert are (6/5) . What is the probability of a customer ordering dessert?
Answer by Edwin McCravy(20086)

Odds in favor = successes:failures = successes/failures

Odds against = failures:successes = failures/successes

probability = successes/(successes + failures)

---------------------------------
odds against a customer ordering dessert are (6/5)
So failures = 6, successes = 5

probability = successes/(successes + failures) = 5/(5+6) = 5/11

Edwin